### v1

@@ Abstract
High-energy neutrinos provide a potentially powerful and distinctive probe for dark matter (DM) - neutrino interactions, particularly in environments with enhanced DM densities, such as the DM spikes predicted to form around supermassive black holes (SMBHs) at the center of active galactic nuclei (AGN). Recent observations by the IceCube Neutrino Observatory, which identified four AGN, namely TXS 0506+056, NGC 1068, PKS 1424+240, and NGC 4151 as neutrino sources, provide a unique opportunity to search for signatures of these interactions. In this study, we use IceCube data to derive the most stringent constraints to date on both the energy-dependent and energy-independent DM-neutrino scattering cross-sections. We perform a statistical analysis using data from individual sources as well as a combined (stacked) analysis of all four sources. Our strongest limits arise from the stacking analysis, yielding an upper bound of $\sigma_{0}\leq 8\times 10^{-39}$ cm2 for an energy-independent cross-section and $\sigma_{0}\leq 10^{-39}$ cm2 for a linearly energy-dependent cross-section, both at 90% confidence level, particularly in scenarios involving the adiabatic growth of black holes.
@@ IV Neutrino flux attenuation due to interaction with DM
where, $\tau=\Sigma_{\chi}(r)/m_{\chi}$ is the accumulated column density over the DM mass, $E_{\nu}$ and $E_{\nu}^{\prime}$ are the outgoing and incoming neutrino energies, respectively. The first term in this equation represents the attenuation of the neutrino flux?, while the second term represents the redistribution of high energy neutrinos in the lower energy bins, due to scattering with DM.
@@ V.1 Likelihood-based $\chi^{2}$ analysis
where $i$ corresponds to the energy bins for which the experimental event number $N_{i}^{ex}\neq 0$ and $j$ corresponds to the energy bins for which $N_{j}^{ex}=0$. The quantities $N_{i}^{th}$ or $N_{j}^{th}$ corresponds to the theoretical estimate of the number of neutrino events from the flux model, after taking into account DM-neutrino scattering and standard neutrino oscillations over astrophysical distances. Here $N^{ex}=n_{s}$ as listed in Table 1. We perform a stacking analysis by combining $\chi^{2}$ from all sources (NGC 1068, TXS 0506+056, PKS 1424+240 and NGC 4151), i.e.
@@ V.2 Results for constant cross-section
    where $n_{s}$ is the total number of neutrino events over all energy bins. From Eq. (19), the 90% upper bound on the DM-$\nu$ cross-section is obtained as
@@
    where, $c\propto\frac{1}{\sqrt{n_{s}}\Sigma_{\chi}}$. The strongest bound observed for NGC 1068 is attributed to the fact that the factor $c$ calculated from the $\chi^{2}$ analysis is the smallest for NGC 1068, due to the observed number of events for this source being the highest. We also find that for more massive SMBHs, the DM density $\rho_{\chi}$ peaks at larger radii, with lower spike height, leading to a smaller column density $\Sigma_{\chi}$ along the neutrino’s propagation to the Earth. Since NGC 1068 hosts the least massive SMBH among the four sources considered, it correspondingly yields the largest $\Sigma_{\chi}$. Consequently, a combination of the largest signal $n_{s}$ and largest column density produces the strongest constraints for this source. As a result, NGC 1068 largely dominates the results of the stacking analysis.
  • For the benchmark models BM2, BM3, BM2 and BM3, where DM-annihilation effects are non-zero, PKS 1424+240 provides the strongest constraint for most of the range of $m_{\chi}$, as shown in Fig. 2. The reason behind this is that DM spikes with higher density get depleted faster due to DM self-annihilation resulting in large decrease in $\Sigma_{\chi}$ as shown for NGC 1068 and NGC 4151 in Fig. 1 and discussed in section III.2. In comparison, the decrease in $\Sigma_{\chi}$ for TXS 0506+056 and PKS 1424+240 at the emission radii of neutrinos (blue shaded region in Fig. 1) is smaller for most $m_{\chi}$ values because of their more massive black holes. Because of a larger $\Sigma_{\chi}$ compared to the NGCs, PKS 1424+240 and TXS 0506+056 give stronger constraint on $\sigma_{0}$ in these benchmark models for most of the $m_{\chi}$ range.
@@ VI Model Interpretation
In this section, we will interpret our results in the parameter space of a well-motivated model of DM with significant coupling to neutrinos. We consider a scenario in which the SM is extended to a gauged and anomaly free $U(1)_{L_{\mu}-L_{\tau}}$ group [64, 65]. Spontaneous symmetry breaking leads to a massive gauge boson $Z^{\prime}$ that couples primarily to $2^{\rm nd}$ and $3^{\rm rd}$ generation SM leptons ($\mu$, $\tau$, $\nu_{\mu}$, $\nu_{\tau}$) and their anti-particles.
@@
with the electromagnetic current for all SM fermions represented by $\mathcal{J}_{EM}^{\mu}=\bar{f}\gamma^{\mu}f$. Although kinetic mixing induces a coupling of all other SM fermions to the $Z^{\prime}$, these interactions are subdominant. We consider a DM $\chi$ charged under this symmetry, with $g_{\chi}\equiv Q_{\mu-\tau}^{\chi}g_{\mu-\tau}$ is the coupling of $Z^{\prime}$ to DM. We assume that DM is vector-like under this gauge symmetry such that the charge ($Q_{\mu-\tau}^{\chi}$) can differ from unity [68, 69]. For illustration, we focus on two benchmark DM scenarios (pseudo-Dirac fermion and complex scalar), in which the corresponding urrent is given by
@@ VI.2 Dark matter - Neutrino Scattering
In Fig. 4, we show the results of our stacking analysis overlaid on the $U(1)_{L_{\mu}-L_{\tau}}$ model parameter space, plotting the gauge coupling $g_{\mu-\tau}$ vs gauge boson mass $m_{Z^{\prime}}$. On the left and right panels we showcase the results of the pseudo-Dirac fermion and complex scalar DM respectively, assuming $m_{Z^{\prime}}=3m_{\chi}$ and $g_{\chi}=1$. The color lines represent our constraints incorporating the different spike density profiles after stacking the results from TXS 0506+056, PKS 1424+240, NGC 4151 and NGC 1068 (the blue dot-dashed line in each panel of Figs. 2 and 3). The solid blue and red lines represent the BM1 and $\rm BM1^{\prime}$ models, respectively. The dashed lines correspond to the BM2 and $\rm BM2^{\prime}$ models, while the dotted lines represent the BM3 and $\rm BM3^{\prime}$ spike density models, respectively.
The gray shaded regions indicate constraints on the model from arising from various experimental and observational searches. The regions associated with the neutrino experiments (i.e. Borexino, KamLand, Super-K) correspond to parameter space in which galactic center DM annihilates into neutrinos; see [69] for further details. Constraints from early-universe cosmology, astrophysical observations and terrestrial probes, including the latest results from the muon g-2 measurements, are discussed in [67].
In this model DM can also interact with electrons or nucleons, via kinetic mixing. However, constraints from direct detection experiments on this space are relatively weak (due to the suppressed kinetic mixing) and are therefore not shown. For brevity, we also omit other existing model constraints that are subdominant compared to those presented here.
@@
Fig. 5 illustrates the impact of our stacking analysis on the DM parameter space, compared to other model independent bounds by recasting the thermal relic density in the $\sigma_{0}$ vs $m_{\chi}$ plane. The black solid and dot-dashed curves correspond to the pseudo-Dirac and complex scalar models respectively, with $m_{Z^{\prime}}=3m_{\chi}$ and $g_{\chi}=1$. For illustration, we present the results in the limit $E_{\nu}\ll m_{Z^{\prime}}/m_{\chi}$ corresponding to the energy dependent scattering cross-section scenario shown in Fig. 3. The color shaded regions indicate to the model-independent bounds as discussed above. Overall, we find that our particle physics model independent stacking analysis provides stringent constraints on the DM parameter space for the chosen benchmark parameters.
@@ VII Discussion & Conclusions
Consistent with the literature, we perform our analysis for both a constant cross-section and a cross-section linearly dependent on the neutrino energy. For the constant cross-section, we find that the strongest constraint is $\sigma_{0}\leq 8\times 10^{-39}$ cm2 at 90% CL for 1 keV DM mass derived from the stacking analysis for BM1 spike density profile. This result from the stacking analysis is dominated by NGC 1068 due to the largest observed number of events and expected large column density for this source. The benchmark model BM1, however, does not take into account effects such as gravitational scattering of DM with surrounding stars or self-annihilation of DM. These effects have been incorporated in other benchmark models BM1, BM2/3 and so on, and we found that our constraints on the DM-neutrino scattering cross-section are weaker for those models. Additionally, for those benchmark models, we found that the most stringent constraints on the DM-neutrino scattering cross-section come from sources with heavier SMBHs, making TXS 0506+056 and PKS 1424+240 particularly significant in this context.
@@
In the energy-dependent cross-section scenario, we find that the most stringent constraint for spike profile BM1 arises from NGC 1068, yielding an upper limit of approximately $\sigma_{0}\leq 10^{-39}$ cm2 at 90% CL. This bound is stronger than that derived in the energy-independent (constant cross-section) case. In contrast, for density profiles with a nonzero $\langle\sigma_{a}v\rangle$ (i.e. BM2, $\rm BM2^{\prime}$, BM3, $\rm BM3^{\prime}$), the strongest constraint instead comes from TXS 0506+056. This can be attributed to the detection of the highest-energy neutrinos from this source (with energies up to $\approx$290 TeV), which enhances experimental sensitivity in the energy-dependent regime. Moreover, the constraints obtained from individual sources are further strengthened when performing a stacking analysis for each assumed DM density profile around SMBHs. Collectively, these results yield the most stringent constraints on the DM-neutrino scattering cross-section, in this scenario, to date.
@@ Appendix A Appendixes
Here we show the event distributions in the energy-dependent cross-section case (see Eq. (15)) for the four sources we considered in our analysis. The values for the constant $A$ in Eq. (15) for these plots are 0.0091 (NGC 1068), 0.0114 (NGC 4151), 0.0056 (TXS 0506+056) and 0.0311 (PKS 1424+240).
Figure 6: Event distributions obtained from the IceCube data (blue-solid) and attenuated up to 90% CL (red-dotdashed).

### v2

@@ Abstract
High-energy neutrinos provide a potentially powerful and distinctive probe for dark matter (DM) - neutrino interactions, particularly in environments with enhanced DM densities, such as the DM spikes predicted to form around supermassive black holes (SMBHs) at the center of active galactic nuclei (AGN). Recent observations by the IceCube Neutrino Observatory, which identified four AGN, namely TXS 0506+056, NGC 1068, PKS 1424+240, and NGC 4151 as neutrino sources, provide a unique opportunity to search for signatures of these interactions. In this study, we use IceCube data to derive the most stringent constraints to date on both the energy-dependent and energy-independent DM-neutrino scattering cross-sections. We perform a statistical analysis using data from individual sources as well as a combined (stacked) analysis of all four sources. Our strongest limits arise from the stacking analysis, yielding an upper bound of $\sigma_{0}\lesssim 8\times 10^{-39}$ cm2 for an energy-independent cross-section and $\sigma_{0}\lesssim 10^{-39}$ cm2 for a linearly energy-dependent cross-section, both at 90% confidence level, particularly in scenarios involving the adiabatic growth of black holes.
@@ IV Neutrino flux attenuation due to interaction with DM
where, $\tau=\Sigma_{\chi}(r)/m_{\chi}$ is the accumulated column density over the DM mass, $E_{\nu}$ and $E_{\nu}^{\prime}$ are the outgoing and incoming neutrino energies, respectively. The first term in this equation represents the attenuation of the neutrino flux, while the second term represents the redistribution of high energy neutrinos in the lower energy bins, due to scattering with DM.
@@ V.1 Likelihood-based $\chi^{2}$ analysis
where $i$ corresponds to the energy bins for which the experimental event number $N_{i}^{ex}\neq 0$ and $j$ corresponds to the energy bins for which $N_{j}^{ex}=0$. The quantities $N_{i}^{th}$ or $N_{j}^{th}$ correspond to the theoretical estimate of the number of neutrino events from the flux model, after taking into account DM-neutrino scattering and standard neutrino oscillations over astrophysical distances. Here $N^{ex}=n_{s}$ as listed in Table 1. We perform a stacking analysis by combining $\chi^{2}$ from all sources (NGC 1068, TXS 0506+056, PKS 1424+240 and NGC 4151), i.e.
@@ V.2 Results for constant cross-section
    where $n_{s}$ is the total number of neutrino events over all energy bins. From Eq. (19), the 90% upper bound on the DM-neutrino cross-section is obtained as
@@
    where, $c\propto\frac{1}{\sqrt{n_{s}}\Sigma_{\chi}}$. The strongest bound observed for NGC 1068 is attributed to the fact that the factor $c$ calculated from the $\chi^{2}$ analysis is the smallest for NGC 1068, due to the observed number of events for this source being the highest. We also find that for more massive SMBHs, the DM density $\rho_{\chi}$ peaks at larger radii, with lower spike height, leading to a smaller column density $\Sigma_{\chi}$ along the neutrino’s propagation to the Earth. Since NGC 1068 hosts the least massive SMBH among the four sources considered, it correspondingly yields the largest $\Sigma_{\chi}$. Consequently, a combination of the largest signal $n_{s}$ and the largest column density produces the strongest constraints for this source. As a result, NGC 1068 largely dominates the results of the stacking analysis.
  • For the benchmark models BM2, BM3, BM2 and BM3, where DM-annihilation effects are non-zero, PKS 1424+240 provides the strongest constraint for most of the range of $m_{\chi}$, as shown in Fig. 2. The reason behind this is that DM spikes with higher density get depleted faster due to DM self-annihilation resulting in large decrease in $\Sigma_{\chi}$ as shown for NGC 1068 and NGC 4151 in Fig. 1 and discussed in section III.2. In comparison, the decrease in $\Sigma_{\chi}$ for TXS 0506+056 and PKS 1424+240 at the emission radii of neutrinos (blue shaded region in Fig. 1) is smaller for most $m_{\chi}$ values because of their more massive black holes. Because of a larger $\Sigma_{\chi}$ compared to the NGCs, PKS 1424+240 and TXS 0506+056 give stronger constraints on $\sigma_{0}$ in these benchmark models for most of the $m_{\chi}$ range.
@@ VI Model Interpretation
In this section, we will interpret our results in the parameter space of a well-motivated model of DM with significant coupling to neutrinos. We consider a scenario in which the SM is extended to a gauged and anomaly-free $U(1)_{L_{\mu}-L_{\tau}}$ group [64, 65]. Spontaneous symmetry breaking leads to a massive gauge boson $Z^{\prime}$ that couples primarily to $2^{\rm nd}$ and $3^{\rm rd}$ generation SM leptons ($\mu$, $\tau$, $\nu_{\mu}$, $\nu_{\tau}$) and their anti-particles.
@@
with the electromagnetic current for all SM fermions represented by $\mathcal{J}_{EM}^{\mu}=\bar{f}\gamma^{\mu}f$. Although kinetic mixing induces a coupling of all other SM fermions to the $Z^{\prime}$, these interactions are subdominant. We consider a DM $\chi$ charged under this symmetry, with $g_{\chi}\equiv Q_{\mu-\tau}^{\chi}g_{\mu-\tau}$ is the coupling of $Z^{\prime}$ to DM. We assume that DM is vector-like under this gauge symmetry such that the charge ($Q_{\mu-\tau}^{\chi}$) can differ from unity [68, 69]. For illustration, we focus on two benchmark DM scenarios (pseudo-Dirac fermion and complex scalar), in which the corresponding current is given by
@@ VI.2 Dark matter - Neutrino Scattering
In Fig. 4, we show the results of our stacking analysis overlaid on the $U(1)_{L_{\mu}-L_{\tau}}$ model parameter space, plotting the gauge coupling $g_{\mu-\tau}$ vs gauge boson mass $m_{Z^{\prime}}$. On the left and right panels we showcase the results of the pseudo-Dirac fermion and complex scalar DM, respectively, assuming $m_{Z^{\prime}}=3m_{\chi}$ and $g_{\chi}=1$. The color lines represent our constraints incorporating the different spike density profiles after stacking the results from TXS 0506+056, PKS 1424+240, NGC 4151 and NGC 1068 (the blue dot-dashed line in each panel of Figs. 2 and 3). The solid blue and red lines represent the BM1 and $\rm BM1^{\prime}$ models, respectively. The dashed lines correspond to the BM2 and $\rm BM2^{\prime}$ models, while the dotted lines represent the BM3 and $\rm BM3^{\prime}$ spike density models, respectively.
The gray shaded regions indicate constraints on the model arising from various experimental and observational searches. The regions associated with the neutrino experiments (i.e. Borexino, KamLand, Super-K) correspond to parameter space in which galactic center DM annihilates into neutrinos; see [69] for further details. Constraints from early-universe cosmology, astrophysical observations and terrestrial probes, including the latest results from the muon g-2 measurements, are discussed in [67].
In this model, DM can also interact with electrons or nucleons, via kinetic mixing. However, constraints from direct detection experiments on this space are relatively weak (due to the suppressed kinetic mixing) and are therefore not shown. For brevity, we also omit other existing model constraints that are subdominant compared to those presented here.
@@
Fig. 5 illustrates the impact of our stacking analysis on the DM parameter space, compared to other model-independent bounds by recasting the thermal relic density in the $\sigma_{0}$ vs $m_{\chi}$ plane. The black solid and dot-dashed curves correspond to the pseudo-Dirac and complex scalar models respectively, with $m_{Z^{\prime}}=3m_{\chi}$ and $g_{\chi}=1$. For illustration, we present the results in the limit $E_{\nu}\ll m_{Z^{\prime}}/m_{\chi}$ corresponding to the energy-dependent scattering cross-section scenario shown in Fig. 3. The color shaded regions indicate the model-independent bounds as discussed above. Overall, we find that our particle physics model-independent stacking analysis provides stringent constraints on the DM parameter space for the chosen benchmark parameters.
@@ VII Discussion & Conclusions
Consistent with the literature, we perform our analysis for both a constant cross-section and a cross-section linearly dependent on the neutrino energy. The strongest constraints observed in our analysis are $\sigma_{0}\lesssim 8\times 10^{-39}$ cm2 and $\sigma_{0}\lesssim 10^{-39}$ cm2 in case of constant and energy-dependent cross sections, respectively, at 90% CL for 1 keV DM mass, which is derived from the stacking analysis for BM1 spike density profile.
For context, it is useful to compare our results with existing bounds in the literature. Previous cosmological analyses based on Milky Way satellite galaxy counts have placed upper limits of $\sigma_{\rm DM-\nu}\lesssim 4\times 10^{-34}\,{\rm cm}^{2}(m_{\rm DM}/{\rm GeV})$ at 95% CL for a constant DM–neutrino scattering cross section, corresponding to relic neutrinos with characteristic energies of order $10^{-4}\,\mathrm{eV}$ [82]. These constraints arise from the impact of DM–neutrino interactions on small-scale structure formation in the early Universe. In contrast, our analysis probes the same interaction using high-energy astrophysical neutrinos detected by IceCube, extending up to TeV energies, and provides the most stringent limits to date in this high-energy regime. For completeness, we also note that a comparatively weaker bound, of order $10^{-22}\,{\rm cm}^{2}\,{\rm GeV}^{-1}$, has been reported using the highest-energy event observed by KM3NeT, where only propagation through the Milky Way dark matter halo was considered [83].
@@
Our results from the stacking analysis, in case of constant cross-section, are dominated by NGC 1068 due to the largest observed number of events and the expected large column density for this source. The benchmark model BM1, however, does not take into account effects such as gravitational scattering of DM with surrounding stars or self-annihilation of DM. These effects have been incorporated in other benchmark models BM1, BM2/3 and so on, and we found that our constraints on the DM-neutrino scattering cross-section are weaker for those models. Additionally, for those benchmark models, we found that the most stringent constraints on the DM-neutrino scattering cross-section come from sources with heavier SMBHs, making TXS 0506+056 and PKS 1424+240 particularly significant in this context.
@@
In the energy-dependent cross-section scenario, the most stringent constraint arises for spike profile BM1 from NGC 1068, which is stronger than that derived in the energy-independent (constant cross-section) case. In contrast, for DM density profiles surrounding SMBHs with a nonzero $\langle\sigma_{a}v\rangle$ ($i.e.$, BM2, $\rm BM2^{\prime}$, BM3, $\rm BM3^{\prime}$), the strongest constraint instead comes from TXS 0506+056. This can be attributed to the detection of the highest-energy neutrinos from this source (with energies up to $\approx$290 TeV), which enhances experimental sensitivity in the energy-dependent regime.
In our analysis of model interpretations, we discussed the anomaly-free $U(1)_{L_{\mu}-L_{\tau}}$ model and presented our results from a stacking analysis conducted in the model’s parameter space. This model considers the scenario of pseudo-Dirac and complex scalar DM, where we established the most stringent constraints on the coupling and mediator mass parameters. A recent analysis of DM–neutrino interactions has examined scalar and fermionic DM coupled through various mediators, considering laboratory, cosmological, and astrophysical constraints, where they excluded many previously accepted scenarios [84]. On the other hand, our study focuses on a well-motivated anomaly-free model, deriving high-energy constraints in its parameter space from IceCube observations at TeV energies.
In conclusion, the constraints obtained in this study on DM-neutrino scattering cross-sections from individual sources are significantly enhanced when we conduct a stacking analysis for each assumed DM density profile around SMBHs. Collectively, these findings provide the most stringent constraints on the DM-neutrino scattering cross-section in this context to date.
@@ Appendix A Appendixes
Here we show the event distributions in the energy-dependent cross-section case (see Eq. (15)) for the four sources we considered in our analysis. The values for the constant $A$ in Eq. (15) for these plots are 0.0091 (NGC 1068), 0.0114 (NGC 4151), 0.0056 (TXS 0506+056) and 0.0311 (PKS 1424+240). These results are obtained keeping the spectral index $\Gamma$ fixed to the values derived from the IceCube public data for each source, as listed in Table 1. However, additionally, we found that adjusting the profiling over $\Gamma$ does not significantly impact our results.
Figure 6: Event distributions obtained from the IceCube data in the standard scenario (blue-solid) and in the DM-neutrino interaction scenario with energy-dependent cross-section at 90% CL (red-dotdashed)